OpenStudy (anonymous):

Which value of x makes the equation true? x-5.8=-4.2 x= ?

OpenStudy (anonymous):

@futureescape @Night-Watcher @texaschic101 @shinalcantara @IMStuck LOL tag party :p

OpenStudy (shinalcantara):

you transpose -5.8 to the right and you'll get the answer

OpenStudy (anonymous):

what does transpose mean?

jhonyy9 (jhonyy9):

x=5,8-4,2 =?

OpenStudy (shinalcantara):

or you can simply put it this way.. add 5.8 to both sides of the equation x-5.8+5.8=-4.2+5.8

OpenStudy (anonymous):

add - 4.2 and 5.8

OpenStudy (anonymous):

got it! 10 thanks every 1 :)